A magnetic field of 1800 A/m produces a magnetic flux of 3 x10-5Wb inan iron bar of cross-section area 0.2 cm2 . Calculate the permeability

Grade:12th pass

1 Answers

Eshan
5 years ago
Permeability of a material is the ratio of flux density created within the material to the magnetising field.

Hence\mu=\dfrac{B}{H}=\dfrac{3\times 10^{-5}/0.2\times 10^{-4}}{1800}=8.3\times 10^{-4}